​(Preferred stockholder expected return​) You own 150 shares of Dalton Resources preferred​ stock, which currently sells for $ 46.55 per share and pays annual dividends of $ 2.75 per share.

a. What is your expected​ return?

b. If you require a return of 6 ​percent, given the current​ price, should you sell or buy more​ stock?

a. Your expected return is ____ percent.​(Round to two decimal​ places.)

b. If you require a return of 6 ​percent, the value of the stock for you is ​$____. ​(Round to the nearest​ cent.) Because the expected rate of return is (less than or greater than) your required rate of return or the intrinsic​ value, or because the current market price is (greater than or less than) $ 45.83 ​, the Dalton Resources preferred stock is (undervalued or overvalued) and you should (buy or sell) the stock.  
Answer #1

a.

Expected Return of Preferred Stock = Annual Dividend/Price

Expected Return = 2.75/46.55

Expected Return = 5.91%

b.

Required Rate = 6%

Price = 2.75/0.06 = $45.83

Price at required return rate is higher than current price, so one should sell.